If f(x)=∫x0(t3+7t2+4)dt
then
f′′(x)=
1
Expert's answer
2020-12-07T16:01:48-0500

By Leibniz's rule, we know that if

"f(x)=\int _ \phi^\psi f(t)dt"

Then "f'(x)=f(
 \psi(x)).\psi'(x)-f(\phi(x)).\phi'(x)........(1)"

Now given that,"f(x)=\int_0^x (t^3+7t^2+4)......(2)"

Comparing (2) with (1) we get,

"\psi(x)=x ,\phi(x)=0,"

"f(t)=t^3+7t^2+4"

Therefore by Leibniz's rule we have,

"f'(x)=f(
 \psi(x)).\psi'(x)-f(\phi(x)).\phi'(x)"

So,"f'(x)=(x^3+7x^2+4).(1)-(0).(0)\n=x^3+7x^2+4"

Therefore, "f''(x)=d\/dx(x^3+7x^2+4)=(3x^2+14x)"

So required answer is "f''(x)=3x^2+14x"